A SCIENTIST WANTS TO KNOW THE NUMBER OF FISH IN A CERTAIN LAKE.SHE CAPTURES 15 FISH,PLACES TAGS ON THEM AND RE-RELEASES THEM INTO THE LAKE. THE NEXT DAY SHE CAPTURES 15 MORE FISH . TWO OF THESE FISH HAVE TAGS. aBOUT HOW MANY FISH ARE IN THE LAKE?
Answer by stanbon(48558)

A SCIENTIST WANTS TO KNOW THE NUMBER OF FISH IN A CERTAIN LAKE.SHE CAPTURES 15 FISH,PLACES TAGS ON THEM AND RE-RELEASES THEM INTO THE LAKE. THE NEXT DAY SHE CAPTURES 15 MORE FISH . TWO OF THESE FISH HAVE TAGS. aBOUT HOW MANY FISH ARE IN THE LAKE?
----------------
Fish in the Lake / # caught = # in sample/ sample fish caught
# fish in Lake/ 15 = 15/2
# fish in Lake = 225/2 = 112.5
Rounding down # of fish in the lake = 112
============
Cheers,
Stan H.